Two small metal spheres are 30.0 cm apart. The spheres have equal amount of negative charges and repel from each other with a force of 40.0 N. What is the charge on each sphere?



1
Expert's answer
2021-04-13T07:02:21-0400

We have,

"d= 30cm = 0.3m"

"F = 40N"

"q_1 = q_2 = -q"

We know,

"F = \dfrac{1}{4\pi \epsilon_0} \dfrac{q_1q_2}{d^2}"


After putting values we get,

"40 = \dfrac{1}{4\pi\epsilon_0} \dfrac{q^2}{0.3\times 0.3}"


"q^2 = 40\times 4\pi\epsilon_0\times 0.3 \times 0.3"

We get,

"q = 0.632\times 10^{-3} C"